A valószínűségszámítás II. gyakorlaton írt dolgozat

1. feladat. 2. feladat 3. feladat 4. feladat. 5. feladat

diákok:

Bóka Zsolt  +  +  ±  +  −  I
Csaba Anikó  +  −  +  +  +  I
Darvas Gábor  +  +  −  −  +  I
Fábián Andrea  +  +  +  +  +  I
Fazekas Zoltán  −  −  ±  −  +  N
Garai Krisztina  −  −  ±  −  −  N
Gombos Gábor  −  +  −  +  +  I
Hadházi István  −  −  ±  −  −  N
Héver Timea  −  +  −  +  −  N
Kis–Tóth Attila  ±  −  ±  −  −  N
Kovács Krisztián  −  +  −  +  +  I
Kovács Marianna  −  +  −  −  −  N
Lukács Attila  −  +  −  −  −  N
Major Imre  +  +  −  −  −  N
Makra Tamás  −  −  −  −  −  N
Malbik Erzsébet  −  −  −  −  +  N
Mészáros Fruzsina  +  +  +  +  −  I
Nagy Erzsébet Bernadett  −  −  −  −  −  N
Nyerges Beáta  +  +  +  +  +  I
Oláh Erika  −  −  −  +  +  N
Orosz Erzsébet  −  +  −  −  −  N
Ottmár Ilona  −  −  −  +  +  N
Ölbei Kinga  +  −  −  +  −  N
Peleskey Eleonóra  −  −  −  +  −  N
Szabó Attila  −  +  −  +  −  N
Szuetta Judit  −  −  +  −  −  N
Tarr Levente  −  +  −  +  −  N
Tóth Krisztina  −  −  −  −  −  N
Vécsei Timea  −  −  −  −  −  N
Veligdán Szilvia  +  −  +  +  −  I
Zákány Irma Ildikó  ±  −  +  −  ±  N